Skip to end of metadata
Go to start of metadata

Work continued at both the sodar and the profiler sites.  At the profiler site, we installed the clutter fence on the wind profiler.  Although I've only had a chance to examine a short period of data afterwards, there did appear to be a significant improvement with the wind turbine being less obvious in the signal plots (the fence was installed 14:30 - 15:30 UTC).   At the sodar site, power was finally installed this afternoon and Eckhard got the sodar working soon afterwards.  It is working very well, reporting winds to the 300 - 400 meter range, and RASS temperature to almost as high.  Gary worked on network connections at both sites, although had limited success with various cell modems and router combinations.  It appears we need to purchase a WiFi hotspot for the profiler site, but may get our NOS USB modem working at the sodar site.   Today the weather was cold, with heavy dense fog all morning, particularly in the valley.